Find the value of y. Round your answer to the nearest tenth.

mc076-1.jpg
1
Expert's answer
2019-04-25T02:27:33-0400

Usually in geometry to find the required value you can apply one of the following or their combination:

  • Pythagorean theorem
  • Sine law
  • Cosine law
  • Projections using one of the trigonometrical functions
  • Different formulas, properties, axioms and theorems.

For instance, consider the following problem: find the value of y - length of an arc:



Calculate the projection of the 'diagonal':


"d=30\text{sin}30^\circ=15,"

calculate the length of an arc:


"y=2\pi d/4=\pi d/2=7.5\pi."